# Exploring Fine-Tuning of Large Audio Language Models for Spoken Language Understanding under Limited Speech data

### v1

@@ Abstract
Index Terms—  large audio language models, spoken language understanding, curriculum learning, cross-lingual transfers
@@ 1 Introduction
Large Audio Language Models (LALMs) such as AudioPaLM [1], SALMONN [2], and Qwen-Audio [3] have recently emerged. By integrating pretrained audio encoders with large language models, they unify acoustic and linguistic representations in an end-to-end manner. This design avoids the limitations of ASR-to-LLM cascades in speech-related tasks, which suffer from transcription errors [4, 5] and loss of prosodic or speaker-related information [6, 7].
@@
In this work, we adopt SLU as a benchmark task to evaluate the practical effectiveness of LALM fine-tuning. SLU is a key task in speech processing and an essential component of conversational agents and voice assistants [13, 14]. At the same time, it also reflects the challenge of limited speech resources: paired speech–label data are expensive to gather and annotate [15], making SLU a representative case of low-resource conditions [16]. Prior studies have proposed several approaches to address this limitation; one such approach is transcription-based training [17], leveraging transcripts that are easier to obtain than speech data and can partially compensate for scarcity.
@@ 3.3 Results on Monolingual SLU
Source (FR) Target de es ko pt vi
Speech Scheme
@@
0% Text T 0.6145 0.6282 0.4319 0.5418 0.0839
2% Curr. T 0.6378 0.6499 0.4450 0.5790 0.1158
5% Curr. T 0.6515 0.6548 0.4522 0.5835 0.1213
10% Curr. T 0.6476 0.6561 0.4609 0.5972 0.1333
25% Curr. T 0.6691 0.6615 0.4612 0.6085 0.1398
50% Curr. T 0.6645 0.6552 0.4564 0.5866 0.1529
100% Curr. T 0.6739 0.6683 0.4726 0.6155 0.1556
@@
0% Curr. T+S 0.6372 0.6630 0.4939 0.6349 0.3317
2% Curr. T+S 0.6537 0.6502 0.4996 0.6342 0.2890
5% Curr. T+S 0.6651 0.6685 0.4918 0.6412 0.3287
10% Curr. T+S 0.6601 0.6696 0.4966 0.6436 0.3265
25% Curr. T+S 0.6785 0.6721 0.5119 0.6501 0.3321
50% Curr. T+S 0.6727 0.6726 0.5126 0.6588 0.3269
100% Curr. T+S 0.6787 0.6804 0.5138 0.6503 0.3351
@@ 3.4 Results in Cross-lingual SLU
Table 3 summarizes the results. Overall, we observe consistent improvements with more source-language speech data, with gains persisting across the full range. This indicates a practical path for low-resource languages: under limited target-language data, combining abundant source-language speech data can substantially improve SLU performance. In the target-language, transcripts alone provide a reasonable baseline, but adding even small amounts of speech data yields further improvements when combined. Vietnamese is a notable exception, likely due to its larger typological gap from French and the relatively limited representation of Vietnamese in the model’s pretraining corpus.

### v2

@@ Abstract
Index Terms—  large audio language models, spoken language understanding, curriculum learning, cross-lingual transfer
@@ 1 Introduction
Large Audio Language Models (LALMs) such as AudioPaLM [15], SALMONN [18], and Qwen-Audio [5] have recently emerged. By integrating pretrained audio encoders with large language models, LALMs unify acoustic and linguistic representations in an end-to-end manner, avoiding the limitations of ASR-to-LLM cascades in speech-related tasks, such as transcription error propagation [6, 17] and the loss of paralinguistic information [16, 21].
@@
In this work, we adopt SLU as a benchmark task to evaluate the practical effectiveness of LALM fine-tuning. SLU is a key task in speech processing and a core component of conversational agents and voice assistants [20, 2], while also representing a low-resource setting where paired speech–label data are expensive to collect and annotate [10, 13]. Prior studies have proposed several approaches to address this limitation; one such approach is transcription-based training [19], leveraging transcripts that are easier to obtain than speech data and can partially compensate for scarcity.
@@ 3.3 Results on Monolingual SLU
Source Speech Target de es ko pt vi
@@
0% T 0.6145 0.6282 0.4319 0.5418 0.0839
2% T 0.6378 0.6499 0.4450 0.5790 0.1158
5% T 0.6515 0.6548 0.4522 0.5835 0.1213
10% T 0.6476 0.6561 0.4609 0.5972 0.1333
25% T 0.6691 0.6615 0.4612 0.6085 0.1398
50% T 0.6645 0.6552 0.4564 0.5866 0.1529
100% T 0.6739 0.6683 0.4726 0.6155 0.1556
@@
0% T+S 0.6372 0.6630 0.4939 0.6349 0.3317
2% T+S 0.6537 0.6502 0.4996 0.6342 0.2890
5% T+S 0.6651 0.6685 0.4918 0.6412 0.3287
10% T+S 0.6601 0.6696 0.4966 0.6436 0.3265
25% T+S 0.6785 0.6721 0.5119 0.6501 0.3321
50% T+S 0.6727 0.6726 0.5126 0.6588 0.3269
100% T+S 0.6787 0.6804 0.5138 0.6503 0.3351
@@ 3.4 Results in Cross-lingual SLU
Table 3 summarizes the results. Overall, we observe consistent improvements with more source-language speech data, with gains persisting across the full range. This indicates a practical path for low-resource languages: under limited target-language data, combining abundant source-language speech data can substantially improve SLU performance. Vietnamese is a notable exception, likely due to its larger typological gap from French and its limited representation in the model’s pretraining corpus. In the target-language, transcripts alone provide a reasonable baseline, but adding even small amounts of speech data yields further improvements when combined.